Please show each step? Thank you! Simpify: -5(y+1)+2(3-5y)
Answer by pedjajov(51)

-5(y+1)+2(3-5y)
:
You have to know that
positive times positive is positive
positive times negative is negative
negative times positive is negative
negative times negative is positive
:
First multiply
(-5)*y+(-5)*1+2*3+2*(-5)*y
-5y -5 +6 -10y
:
Now combine those that have just variable y: -5y-10y=-15y
-15y-5+6
:
Now combine numbers: -5+6=1
-15y+1
:
This is your simplified expression